OpenStudy (anonymous):

Solve x^4 = 36x^2. Identify any double roots. A. {0 (double root), 6, –6} B. {–3, 3, 2 (double root)} C. {0, 3, 2} D. {6, –6}

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

\[ x^4 = 36x^2\] \[x^4-36x^2=0\] \[x^2(x^2-36)=0\] is a good start

OpenStudy (anonymous):

you get \[x^2=0\] which means \(0\) is a double root i bet you can find the other two

OpenStudy (anonymous):

Is the correct answer A?

OpenStudy (anonymous):

yep then \[x^2 - 36 = 0 \] \[x^2 = 36\] \[x = 6\] \[x = -6 \]

OpenStudy (anonymous):

so yeah A would be correct :)

OpenStudy (anonymous):

Thank you so much!:)

OpenStudy (anonymous):

No problem also thank @satellite73 :)